# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ Appendix D Omitted Results and Proofs in Section 4
$\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]^{2}\preccurlyeq$ ${\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right).$
And by Lemma E.3,
$\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]^{\intercal}$
$\qquad=\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}{\bm{C}}^{-T}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\intercal}}\right)$
$\qquad=\left\|{{\bm{\phi}}(s^{\prime})}\right\|^{2}\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}{\bm{C}}^{-T}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$
$\qquad\preccurlyeq\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$
$\qquad\preccurlyeq\gamma{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right).$
@@
${\bm{A}}{\bm{A}}^{\intercal}\preccurlyeq$ $2(1+\gamma){\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right),$
after taking expectation
${\mathbb{E}}\left[{{\bm{A}}{\bm{A}}^{\intercal}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (35)
Put everything together (Eqn. (33), Eqn. (34), Eqn. (35)), we have
${\bm{B}}^{2}\preccurlyeq$ $2(1+\beta)\left({{\bm{A}}^{\intercal}{\bm{A}}+{\bm{A}}{\bm{A}}^{\intercal}}\right)+(1+\beta^{-1})\alpha^{2}\left({{\bm{A}}^{\intercal}{\bm{A}}}\right)^{2}$
@@
D.4 Proof of Theorem 4.1
Proof.
Combining Lemma 5.1, Lemma 5.2 and Lemma 5.3 with [Theorem 1 Samsonov et al., 2024b], we have
${\mathbb{E}}^{1/2}\left[{\left({{\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)}\right)^{2}}\right]\lesssim$ $\frac{1}{\sqrt{K(1-\gamma)^{4}\lambda_{\min}}}\Bigg{[}\sqrt{\frac{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}{T}}\left({1+\frac{C_{A}\sqrt{\alpha}}{\sqrt{a}}}\right)+\frac{\sqrt{\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)}}{\sqrt{\alpha a}T}$
$+(1-\alpha a)^{T/2}\left({\frac{1}{\alpha T}+\frac{C_{A}}{\sqrt{\alpha a}T}}\right)\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|\Bigg{]}$
$\lesssim$ $\frac{1}{\sqrt{K(1-\gamma)^{4}\lambda_{\min}}}\Bigg{[}\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+\sqrt{K}(1-\gamma)}{\sqrt{T}}\left({1+\sqrt{\frac{\alpha}{(1-\gamma)\lambda_{\min}}}}\right)$
$+\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+\sqrt{K}(1-\gamma)}{\sqrt{\alpha(1-\gamma)\lambda_{\min}}T}$
$+(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}\left({\frac{1}{\alpha T}+\frac{1}{\sqrt{\alpha(1-\gamma)\lambda_{\min}}T}}\right)\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|\Bigg{]}$
$\lesssim$ $\frac{1}{\sqrt{T}}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\sqrt{\frac{\alpha}{(1-\gamma)\lambda_{\min}}}}\right)$
$+\frac{1}{T}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
$+\frac{1}{T}\frac{(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}}{\sqrt{\alpha}(1-\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}+\frac{1}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|.$
D.5 $L^{p}$ Convergence
Theorem D.1 ($L^{p}$ Convergence).
For any $K\geq(1-\gamma)^{-1}$, $p>2$, $T\geq 2$, $\alpha\in(0,(1-\sqrt{\gamma})/[38(p+\log T)])$ and initialization ${\bm{w}}_{0}\in{\mathbb{R}}^{dK}$, it holds that
${\mathbb{E}}^{1/p}\left[{\left({{\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)}\right)^{p}}\right]\lesssim$ $\frac{\sqrt{p}}{\sqrt{T}}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\frac{\sqrt{\alpha p}+\alpha p}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)$
@@ D.5 $L^{p}$ Convergence
Combining Lemma 5.1, Lemma 5.2 and Lemma 5.3 with [Theorem 2 Samsonov et al., 2024b], we have
${\mathbb{E}}^{1/p}\left[{\left({{\mathcal{L}}\left({\bar{\bm{w}}_{T}}\right)}\right)^{p}}\right]$
@@
D.6 Convergence Results for SSGD with the PMF Representation
In this section, we present without proof, the counterparts of Lemma 5.1, Lemma 5.2 and Lemma 5.3 for SSGD with the PMF representation. These results will additionally depend on $K$. The additional $K$-dependent terms arise because the condition number of ${\bm{C}}^{\intercal}{\bm{C}}$ scales with $K^{2}$ (Lemma E.2). These terms are inevitable within our theoretical framework. By [Theorem 1 Samsonov et al., 2024b], we find that we cannot obtain a convergence rate independent of $K$ for the vanilla form Eqn. (23) without a preconditioner as in our Linear-CTD. A natural question is whether the convergence rate of this algorithm truly depends on $K$. We leave this as future work.
The proofs of these results only require minor modifications to the original proofs. For the sake of simplicity, we will not introduce new notations. Instead, we will only add the subscript $\operatorname{PMF}$ to the original notations to indicate the difference.
According to Eqn. (23) in Appendix 7, the algorithm corresponds to the following linear system for ${\bm{w}}\in{\mathbb{R}}^{dK}$
$\bar{{\bm{A}}}_{\operatorname{PMF}}{\bm{w}}=\bar{{\bm{b}}}_{\operatorname{PMF}},$
@@ D.6 Convergence Results for SSGD with the PMF Representation
$\bar{{\bm{A}}}_{\operatorname{PMF}}$ $=\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}\tilde{{\bm{G}}}(r)}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right],$
$\bar{{\bm{b}}}_{\operatorname{PMF}}=\frac{1}{K+1}{\mathbb{E}}_{s,r}\left\{{\left[{{\bm{C}}^{\intercal}{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)}\right]\otimes{\bm{\phi}}(s)}\right\}.$
Lemma D.1.
For any ${\bm{w}}\in{\mathbb{R}}^{dK}$, it holds that
$\left({{\mathcal{L}}({\bm{w}})}\right)^{2}\leq\frac{16}{K(1-\gamma)^{4}\lambda_{\min}}\left\|{\bar{{\bm{A}}}_{\operatorname{PMF}}\left({{{\bm{w}}}-{\bm{w}}^{\star}}\right)}\right\|^{2}.$
Lemma D.2.
$C_{A,\operatorname{PMF}}\leq 4K^{2},$
$C_{e,\operatorname{PMF}}\leq 4K^{\frac{3}{2}}\left({\sqrt{K}\left\|{{\bm{w}}^{\star}}\right\|+\left({1-\gamma}\right)}\right),$
$\operatorname{tr}\left({{\bm{\Sigma}}_{e,\operatorname{PMF}}}\right)\leq 18K^{2}\left({\left\|{{\bm{w}}^{\star}}\right\|_{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}^{2}+K(1-\gamma)^{2}}\right).$
Lemma D.3.
For any $p\geq 2$, let $a=(1-\sqrt{\gamma})\lambda_{\min}/2$ and $\alpha_{p,\infty}=(1-\sqrt{\gamma})/(38K^{2})$ ($\alpha_{p,\infty}p\leq 1/2$). Then for any $\alpha\in\left({0,\alpha_{p,\infty}}\right)$, ${\bm{u}}\in{\mathbb{R}}^{dK}$ and $n\in{\mathbb{N}}$
${\mathbb{E}}^{1/p}\left[{\left\|{{\bm{\Gamma}}_{1:n,\operatorname{PMF}}^{(\alpha)}{\bm{u}}}\right\|^{p}}\right]\leq\left({1-\alpha a}\right)^{n}\left\|{{\bm{u}}}\right\|.$
Appendix E Other Technical Lemmas

### v2

@@ Appendix E Omitted Results and Proofs in Section 4
$\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right]^{2}\preccurlyeq$ ${\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right).$
And by Lemma G.3,
$\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]^{\top}$
$\qquad=\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}{\bm{C}}^{-T}\tilde{{\bm{G}}}^{\top}(r){\bm{C}}^{\top}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\top}}\right)$
$\qquad=\left\|{{\bm{\phi}}(s^{\prime})}\right\|^{2}\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}{\bm{C}}^{-T}\tilde{{\bm{G}}}^{\top}(r){\bm{C}}^{\top}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)$
$\qquad\preccurlyeq\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)$
$\qquad\preccurlyeq\gamma{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right).$
@@
${\bm{A}}{\bm{A}}^{\top}\preccurlyeq$ $2(1+\gamma){\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right),$
and after taking expectation
${\mathbb{E}}\left[{{\bm{A}}{\bm{A}}^{\top}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (38)
By putting everything together (Eqn. (36), Eqn. (37), Eqn. (38)), we have
${\bm{B}}^{2}\preccurlyeq$ $2(1+\beta)\left({{\bm{A}}^{\top}{\bm{A}}+{\bm{A}}{\bm{A}}^{\top}}\right)+(1+\beta^{-1})\alpha^{2}\left({{\bm{A}}^{\top}{\bm{A}}}\right)^{2}$
@@
E.4 $L^{p}$ Convergence
Theorem E.1 ($L^{p}$ Convergence).
For any $K\geq(1-\gamma)^{-1}$, $p>2$ and $\alpha\in(0,(1-\sqrt{\gamma})/[38(p+\log T)])$, it holds that
${\mathbb{E}}^{1/p}\left[{\left({{\mathcal{L}}\left({\bar{\bm{\theta}}_{T}}\right)}\right)^{p}}\right]\lesssim$ $\frac{\sqrt{p}}{\sqrt{T}}\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\frac{\sqrt{\alpha p}+\alpha p}{\sqrt{(1-\gamma)\lambda_{\min}}}}\right)$
@@ E.4 $L^{p}$ Convergence
Combining Lemma 5.1, Lemma 5.2 and Lemma 5.3 with [Theorem 2 54], we have
${\mathbb{E}}^{1/p}\left[{\left({{\mathcal{L}}\left({\bar{\bm{\theta}}_{T}}\right)}\right)^{p}}\right]$
@@
E.5 Convergence Results for SSGD with the PMF Representation
In this section, we present the counterparts of Lemma 5.1, Lemma 5.2, Lemma 5.3 and Theorem 4.1 for stochastic semi-gradient descent (SSGD) with the probability mass function (PMF) representation. These results will additionally depend on $K$. The additional $K$-dependent terms arise because the condition number of ${\bm{C}}^{\top}{\bm{C}}$ scales with $K^{2}$ (Lemma G.2). These terms are inevitable within our theoretical framework. The proofs of these results require only minor modifications to the original proofs, and we omit them for brevity.
In fact, in Appendix F, we validate some theoretical results through numerical experiments. To be concrete, we find that empirically, as $K$ increases, to ensure convergence, the step size of the vanilla algorithm in [4, Section 9.6] indeed needs to decay at a rate of $K^{-2}$. In contrast, the step size of our Linear-CTD does not need to be adjusted when $K$ increases. Moreover, we find that Linear-CTD empirically consistently outperforms the vanilla algorithm under different $K$.
Recall Eqn. (26), the updating scheme of the algorithm is
${\bm{W}}_{t}\leftarrow$ ${\bm{W}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left({{\bm{p}}_{{\bm{w}}_{t-1}}(s_{t})-{\bm{p}}_{\bm{\mathcal{T}}_{t}^{\pi}{\bm{\eta}_{{\bm{w}}_{t-1}}}}(s_{t})}\right)^{\top}{\bm{C}}^{\top}{\bm{C}}$
$=$ ${\bm{W}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{W}}_{t-1}-{\bm{\phi}}(s_{t+1})^{\top}{\bm{W}}_{t-1}\tilde{{\bm{G}}}^{\top}(r_{t})-\frac{1}{K+1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t})-\bm{1}_{K}}\right)^{\top}}\right]{\bm{C}}^{\top}{\bm{C}},$
which is equivalent to
${\bm{W}}_{t}{\bm{C}}^{\top}{\leftarrow}$ ${\bm{W}}_{t{-}1}{\bm{C}}^{\top}{-}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{W}}_{t{-}1}{\bm{C}}^{\top}{-}{\bm{\phi}}(s_{t{+}1})^{\top}{\bm{W}}_{t{-}1}{\bm{C}}^{\top}({\bm{C}}\tilde{{\bm{G}}}(r_{t}){\bm{C}}^{-1})^{\top}-\frac{1}{K{+}1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}){-}\bm{1}_{K}}\right)^{\top}{\bm{C}}^{\top}}\right]{\bm{C}}{\bm{C}}^{\top},$
here we drop the additional $2\iota_{K}$ in the step size for brevity. Letting ${\bm{\Theta}}_{\operatorname{PMF},t}:={\bm{W}}_{t}{\bm{C}}^{\top}$ be the CDF parameter, the algorithm becomes
${\bm{\Theta}}_{\operatorname{PMF},t}{\leftarrow}$ ${\bm{\Theta}}_{\operatorname{PMF},t-1}{-}\alpha{\bm{\phi}}(s_{t})\left[{{\bm{\phi}}(s_{t})^{\top}{\bm{\Theta}}_{\operatorname{PMF},t{-}1}-{\bm{\phi}}(s_{t{+}1})^{\top}{\bm{\Theta}}_{\operatorname{PMF},t{-}1}({\bm{C}}\tilde{{\bm{G}}}(r_{t}){\bm{C}}^{-1})^{\top}{-}\frac{1}{K{+}1}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}){-}\bm{1}_{K}}\right)^{\top}{\bm{C}}^{\top}}\right]{\bm{C}}{\bm{C}}^{\top}.$ (39)
Here, we add the subscript $\operatorname{PMF}$ to the original notations to indicate the difference. Then, the algorithm corresponds to the following linear system for ${\bm{\theta}}\in{\mathbb{R}}^{dK}$
$\bar{{\bm{A}}}_{\operatorname{PMF}}{\bm{\theta}}=\bar{{\bm{b}}}_{\operatorname{PMF}},$
@@ E.5 Convergence Results for SSGD with the PMF Representation
$\bar{{\bm{A}}}_{\operatorname{PMF}}$ $=\left[{\left({{\bm{C}}{\bm{C}}^{\top}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}{\bm{C}}^{\top}({\bm{C}}\tilde{{\bm{G}}}(r_{t}){\bm{C}}^{-1})}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right],$
$\bar{{\bm{b}}}_{\operatorname{PMF}}=\frac{1}{K+1}{\mathbb{E}}_{s,r}\left\{{\left[{{\bm{C}}{\bm{C}}^{\top}{\bm{C}}\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)}\right]\otimes{\bm{\phi}}(s)}\right\}.$
Compared to our Linear-CTD (Eqn. (16)), this algorithm has an additional matrix ${\bm{C}}{\bm{C}}^{\top}$ with the condition number of order $K^{2}$ (see Lemma G.2).
Now, we are ready to state the theoretical results for the algorithm.